OpenStudy (anonymous):

write the equation of each line in the given form the line through (0,-2) and (4,6) in point slope form

OpenStudy (anonymous):

point slope form is y-y sub 1=m(x-x sub1) to find the slope you do m= (y sub1 - y sub2)/(x sub 1 - x sub 2)

OpenStudy (anonymous):

so the slope would be (-2+6)/(0-4) which equals 4/-4

OpenStudy (anonymous):

so the point slope form would be y+2=-1(x-0) first you distribute the -1: y+2=-1x then you subtract the 2 from both sides: y=-x-2